OpenStudy (anonymous):

An athlete twirls a 16 kg hammer tied to the end of a 1.3-m rope horizontally about his head. The hammer moves at 40.84 m/s. What is the force of tension in the rope?

OpenStudy (anonymous):

the net force on the is mv^2/r so Fnet = mv^2/r Ft = mv^2/r Ft = (16)(40.84)^2/(1.3m)